Johnson Company manufactures one of the components, BELTO, required for its main product, ALTO. The manufacturing costs per unit for next month’s production of 10,000 units of BELTO when manufactured in house are:

             

                              Direct materials                                                                 $20

                              Direct labor                                                                         15

                              Variable factory overhead                                                   16

                              Allocated fixed factory overhead                                        24

                              Total costs                                                                         $75

             

Johnson’s fixed factory overhead costs are not controllable and cannot be avoided in the short run. Assume that, for its next month’s requirement, Johnson can buy all 10,000 units of BELTO from another manufacturer at a price of $64 per unit. In such a situation, the machines currently used to make BELTO could be temporarily used to generate additional monthly contribution of $150,000 to temporarily meet the additional demand for one of Johnson’s regular products. Under this situation, if Johnson buys BELTO from outside

Johnson’s monthly profit will increase by $20,000

Johnson’s monthly profit will increase by $40,000

Johnson’s monthly profit will increase by $490,000

Johnson’s monthly profit will decrease by $40,000
Answer #1

Relevant cost to Make

Variable Cost Per Unit = Direct Material +Direct labor +Variable Factory overheads = 20+15+16 = $51

Cost = 10,000*51 = 510,000

Relevant cost to Buy

Cost per Unit = 64

Cost = 64*10,000-150,000

Addition income generated = 150,000

total Relevant cost to buy=640,000-150,000= 490,00

Financial Advantage of Buying the Prodcut = 510,000-490,000 = $20,000

Answer : Johnson's monthly Profit will increase by $ 20,000